Car parking charges.
A local council raises the price of car parking from E3 per day to {5 per day and finds that usage of car parks contracts from 1,200 cars a day to 900 cars per day.
Calculate the price elasticity of demand for this price change and calculate whether total revenue from the car park rises or falls.
1
Expert's answer
2020-10-15T02:53:15-0400
"p_1=3, p_2=5"


"Q_1=1200,Q_2=900"


"E=\frac {Q_1-Q_2}{p_1-p_2}\times \frac{p_1}{Q_1}=-0.375"


"TR_1=p_1Q_1=3600"

"TR_2=p_2Q_2=4500"


"\varDelta TR=4500-3600=900"